The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu, Dominik J. Bomans, and Eva K. Grebel

   mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/mcnews/MCNews.html
http://www.astro.uni-wuerzburg.de/~grebel/MCNews.html

No. 8 Dec 16, 1996

Table of Contents

News and Views...............1 Paper Abstracts...............2 Job Advertisements..........6 Meeting Announcements..7

.            +           .           .         .
.      +    .       .     *    +              .
.     .     *     .   /*\      .     *       .    .
*      ,    . .    /***\   .     .      .    *
.    .    .   .  +   <*******>    +      .    .      .
.     +      .      /*****\   .      .    *      +
+  .  .    *        /*********\    *    +        .   ,
.      .   .   <*************>   .      +      .
.      +     .     .      #   .        .      *       .
____________________________#______________________________
Happy Holidays and Best Wishes for the New Year!
___________________________________________________________

In Memorial

We regret to inform our readers that Dr. V. Ambartsumian passed away on August 12, 1996. A memorial article about Dr. Arbartsumian and his contributions to astrophysics was written by Dr. G.

Israelian and can be downloaded from: http://www.astro.uiuc.edu/mcnews/ambartsumian.ps.gz

Paper Abstracts

Interactions of high velocity clouds with the disk of the LMC

M. Braun^{1}

1 Astrophysikalisches Institut Potsdam, An der Sternwarte 16, D-14482 Potsdam, Germany

High-resolution (∼1.′0-2.′4) IRAS maps, optical broad-band, and narrow-band images as well as H i λ 21 cm data have been used in order to analyse the environment of two large-scale distortions (LMC 4 and DEM 268) in the disk of the Large Magellanic Cloud. This study has revealed cone-like structures connected to DEM 268 and LMC 4 indicating high-velocity cloud impacts as the most probable source of the formation of these two distortions. In both cases, the tip of the cone, which is a place of increased star formation, has a shape which is typical of supersonic motion of gaseous matter through a compressible medium. Expansion velocities of the cone-like structures of about 12 km·s^{−1} can be derived from the H i λ 21 cm data. From these velocities, ages of the distortions of about 40 Myr (LMC 4) and 36 Myr (DEM 268) are estimated. The energy input for the formation of the distortions might have been provided by high velocity clouds with masses of the order of 10^{5} M_{⊙} (assuming an initial velocity of 100 km·s^{−1} relative to the disk of the LMC). Accepted by: Astronomische Nachrichten (AN 317, 369) For preprints, contact MBraun@aip.de

Also available by anonymous ftp at ftp.aip.de/incoming/mbraun/machcone.ps.gz
or from the URL http://www.aip.de:8080/science/preprints/mbr001.html

Star Formation Activity in the Large Magellanic Cloud: FIR Emission From IRAS High Resolution Data

Douglas A. Caldwell^{1,2} and Marc L. Kutner^{1,2}

1 Department of Physics, Applied Physics and Astronomy, Rensselaer Polytechnic Institute, Troy, New York 12180, USA 2 National Radio Astronomy Observatory, 949 N. Cherry Ave, Campus Bldg 65, Tucson, AZ

We present an investigation of the properties of 21 molecular clouds in the LMC. Our data consist of 60 μm and 100 μm IRAS images which we compare with fully-sampled ^{12}CO (J=1→0) maps. The CO data were taken on the SEST 15m which has a beam-width at 2.6 mm, of 45”, or 10 pc at the distance of the LMC. We use the IRAS high resolution reprocessed data, with approximate resolutions of 60”, and 75” at 60 μm and 100 μm respectively. The clouds mapped are in three regions: two complexes south of 30 Doradus, and one at the HII region N11. We measure the far–infrared luminosities for each cloud which has a good CO–FIR correspondence, and compare these results with those from similar studies done on Milky Way molecular clouds. The far-infrared luminosities range from 3.5 × 10^{4} to 2.8 × 10^{6} L_{⊙}, with an average value of 4 × 10^{5} L_{⊙}. This average is lower than that seen for HII regions in the inner Milky Way (Scoville & Good 1989, ApJ 339,149), and slightly higher than that for outer Galaxy clouds (Mead et al. 1990, ApJ 354, 492). Virial masses of these clouds are from 4 × 10^{4} to 1 × 10^{6} M_{⊙}. Star formation activity (L_{FIR}/M_{VIR}), defined as the ratio of the luminosity from recently formed stars to the total cloud mass, has a two order of magnitude range for any M_{VIR}. In addition, we find that L_{FIR}/M_{VIR} is independent of cloud mass over the entire range of measured virial masses. These results indicate that although less luminous in CO and FIR than their inner Milky Way counterparts, LMC molecular clouds are undergoing significant massive star formation.

L_{FIR} is used to determine the numbers and types of stars embedded in the GMCs. We compare these findings with those from the Milky Way and discuss the implications for star formation theories. Accepted by: Astrophysical Journal, to appear Dec 1, 1996

For preprints, contact caldwell@charon.phys.rpi.edu
or from the URL http://www.rpi.edu/~caldwd/work.htmd

The asymmetric radio remnant of SN 1987A

B. M. Gaensler^{1,2}, R. N. Manchester^{1}, L. Staveley-Smith^{1}, A. K. Tzioumis^{1}, J. E. Reynolds^{1} and M. J. Kesteven^{1}

1 Australia Telescope National Facility, CSIRO, PO Box 76, Epping, NSW 2121, Australia 2 Astrophysics Department, School of Physics A29, University of Sydney, NSW 2006, Australia

We present seven years of radio observations of SN 1987A made with the Australia Telescope Compact Array. At 1.4, 2.4, 4.8 and 8.6 GHz, the flux density of the radio remnant has increased monotonically since emission was redetected 1200 days after the explosion. On day 3200, the remnant was expanding at 2800 ± 400 km s^{−1}, which we interpret as indicating significant deceleration of the fastest moving ejecta. Since day 1787 the spectral index has remained constant at α = −0.95 ± 0.04 (S ∝ ν^{α}). These observations are all consistent with the shock having encountered a denser, shocked, component of the progenitor’s stellar wind. At the current rate of expansion, the shock is expected to encounter the inner optical ring in the year 2006±3, in line with predictions made by hydrodynamic simulations.

Using super-resolution, we have also obtained 9 GHz images of the remnant (resolution ≈ 0′′.5) at four epochs. The emission is distributed around the rim of a near-circular shell, but has become increasingly asymmetric with time. There are two “hotspots” to the east and west, aligned along the major axis of the optical ring. This morphology is most likely indicative of an axisymmetric circumstellar medium into which the shock is expanding, consistent with present understanding of the progenitor star and its environment. The two hotspots are increasing in flux density at different rates, which may indicate directional anisotropies in the ejecta. We believe that the northern and southern regions of the remnant are encountering a shocked wind which is less dense and also further from the progenitor star than that in the bright regions of emission. As a result, these regions should eventually brighten and/or extend. Accepted by: The Astrophysical Journal

For preprints, contact b.gaensler@physics.usyd.edu.au
Also available from the URL http://www.physics.usyd.edu.au/~bmg/sn1987a/

On the evolutionary phase and mass loss of the Wolf-Rayet-like stars in R136a

Alex de Koter^{1}, Sara R. Heap^{2}, Ivan Hubeny^{3}

1 Advanced Computer Concepts, Code 681, NASA/Goddard Space Flight Center, Greenbelt MD 20771 2 Laboratory for Astronomy and Solar Physics, Code 681, NASA/Goddard Space Flight Center, Greenbelt MD 20771 3 Universities Space Research Association, Code 681, NASA/Goddard Space Flight Center, Greenbelt MD 20771

We report on a systematic study of the most massive stars, in which we analyzed the spectra of four very luminous stars in the Large Magellanic Cloud. The stars lie in the 30 Doradus complex, three of which are located in the core of the compact cluster, R136a (R136a1, R136a3, and R136a5), and the fourth (Melnick 42), located about 8” north of R136a. Low-resolution spectra (< 200 km·s^{−1}) of these four stars were obtained with the GHRS and FOS spectrographs on the Hubble Space Telescope. The GHRS spectra cover the spectral range from 1200 to 1750 Å, and the FOS spectra from 3200 to 6700 Å. We derived the fundamental parameters of these stars by fitting the observations by model spectra calculated with the isa-wind code of de Koter et al. We find that all four stars are very hot (∼ 45 kK), luminous, and rich in hydrogen. Their positions on the HR-diagram imply that they are stars with masses in the range 60-90 M_{⊙} that are 2 million years old at most, and hence, they are O-type main-sequence stars still in the core H-burning phase of evolution. Nevertheless, the spectra of two of the stars (R136a1, R136a3) mimic those of Wolf-Rayet stars in showing very strong He ii emission lines. According to our calculations, this emission is a natural consequence of a very high mass-loss rate.

We conjecture that the most massive stars in R136a – those with initial masses of ∼ 100 M_{⊙} or more – are born as WR-like stars and that the high mass loss may perhaps be connected to the actual stellar formation process. Because the observed mass-loss rates are up to three times higher than assumed by evolutionary models, the main sequence and post-main sequence tracks of these stars will be qualitatively different from current models. The mass-loss rate is 3.5 to 8 times that predicted by the analytical solutions for radiation-driven winds of Kudritzki et al. (1989). However, using sophisticated Monte-Carlo calculations of radiative driving in unified model atmospheres, we show that – while we cannot say for sure what initiates the wind – radiation pressure is probably sufficient to accelerate the wind to its observed terminal velocity, if one accounts for the effects of multiple photon scattering in the dense winds of the investigated stars. Accepted by the Astrophysical Journal

For preprints, contact alex@homie.gsfc.nasa.gov
or by anonymous ftp to homie.gsfc.nasa.gov cd pub

A photometric and spectroscopic study of the association LH 47 in the superbubble N 44 in the LMC

J.-M. Will^{1}, D.J. Bomans^{2}, A. Dieball^{1}

1 Sternwarte, Univ Bonn, D-53121 Bonn, Germany 2 Deptm. of Astronomy, Univ. Illinois at Urbana Champaign, Il 61081, USA

The OB association LH 47 is analyzed with CCD photometry. For the bright blue stars spectra in the optical and UV have been taken. The observed U, B, and V magnitudes and colors are converted to T_{eff} and L. From isochrone fits to the color magnitude diagram we find an age of the association of about 5 to 6 Myr. The stellar population outside the shell is of similar age, so there is no direct evidence for sequential star formation.

The slope of the luminosity function (γ = 0.24) agrees with what has been found for other young associations. Two methods to derive the IMF give basically the same result: x = 1.37 (counting stars along the main sequence) and x = 1.22 (counting stars between evolutionary tracks). The IMF of the stars outside the shell is slightly steeper. The values found are very similar to those for other associations, supporting the idea of a universal IMF. LH 47 turns out to be a well behaved young association embedded in a molecular cloud. Accepted by: Astronomy & Astrophysics Supplements

For preprints, contact bomans@astro.uiuc.edu or JEAN-MARIE WILL@HP-Germany-om5.om.hp.com
Also available from the URL http://www.astro.uiuc.edu/~bomans/preprints.html
or http://www.astro.uni-bonn.de/~jmwill/publications.html

Properties of LMC Planetary Nebulae and Parent Populations in the MACHO Database

D.R. Alves^{1} C. Alcock^{2}, R.A. Allsman^{3}, T.S Axelrod^{4}, A.C Becker^{6}, D.P. Bennett^{7}, K.H. Cook^{2}, K.C. Freeman^{5}, K. Griest^{8}, J. Guern^{8}, M.J. Lehner^{8}, S.L. Marshall^{2},

D. Minniti^{2}, B.A. Peterson^{5}, P.J. Quinn^{9}, M.R. Pratt^{6}, A.W. Rodgers^{5}, C.W. Stubbs^{6}, W. Sutherland^{10}, D. Welch^{11}, (MACHO Collaboration)

1 UC Davis/LLNL, ^{2} LLNL, ^{3} ANUSF, ^{4} MSSSO/LLNL, ^{5} MSSSO, ^{6} CfPA/U Wash., 7 Notre Dame, ^{8} CfPA/UCSD, ^{9} ESO, ^{10} Oxford, ^{11} McMaster

The MACHO microlensing experiment’s time-sampled photometry database contains blue and red lightcurves for nearly 9 million stars in the central bar region of the Large Magellanic Cloud (LMC). We have identified known LMC Planetary Nebulae (PN) in the database and find one, Jacoby 5, to be variable. We additionally present data on the “parent populations” of LMC PN, and discuss the star formation history of the LMC bar. To appear in: Proceedings of IAU Symposium 180, Planetary Nebulae For preprints, contact alves@igpp.llnl.gov

or from the URL http://wwwmacho.mcmaster.ca

The HI Supergiant Shells in the Large Magellanic Cloud

Sungeun Kim^{1} and Lister Staveley-Smith^{2}

1 Mount Stromlo and Siding Spring Observatories 2 Australian Telescope National Facility

We present the results of an HI aperture synthesis mosaic of the Large Magellanic Cloud (LMC), made recently with the Australian Telescope Compact Array (ATCA). The resolution of the mosaicked images is 1.′0 (15 pc, using a distance to the LMC of 50 kpc). In contrast with its appearance at other wavelengths, the LMC is remarkably symmetric in HI on the largest scales, with the bulk of the HI residing in a disk of diameter 8.^{∘}4 (7.3 kpc). Outer spiral structure is clearly seen, though the features appear to be caused by differential rotation, therefore transient in nature. On small to medium scales, the combined action of numerous shells and supergiant shells dominates the structures and motions of the HI in the LMC. Some supergiant shells show good correlation between HI and Hα structures. We compare the results with a new wide-field Hα image. To appear in: ”Star Formation Near and Far”, Proceedings of the 7th annual Astrophysics Conference in Maryland For preprints, contact sek@mso.anu.edu.au

Job Advertisements

Postdoctoral Positions at UNAM

Applications are invited for several postdoctoral positions at the Instituto de Astronomia at the Universidad Nacional Autonoma de Mexico (UNAM), in Mexico City, with branches in Ensenada, Baja California and Morelia, Michoacan.

Positions are available in theory and observation with emphasis on the following research areas and places: Star Formation (Morelia), Galactic and Extragalactic Dynamics (Ensenada), IR Studies of Galaxies, Blue Galaxies and Extragalactic HII Regions, Interacting Pairs of Galaxies, Induced Starbursts and Active Galactic Nuclei, Ionized Gas in Elliptical and Lenticular Galaxies, Chemical Evolution of Galaxies and Physics of Diffuse Nebulae (Mexico City).

Successful applicants will have access to the National Observatory of Mexico in San Pedro Martir, Baja California, with 0.84m, 1.5m and 2.1m telescopes; low and high resolution spectrographs, optical CCD cameras, an IR camera/spectrograph, a mid-IR camera, a Fabry–Perot scanning Interferometer and Stromgren and Johnson photometers. A network of workstations is available with access to a CRAY-Y/MP (Mexico City) and a SG mini-supercomputer (Ensenada). Information can be found on

the WEB at http://www.astroscu.unam.mx/, or requested from aguilar@bufadora.astrosen.unam.mx

Positions are for 1 year with option of renewal for another. Applicants should send a letter of intention, curriculum vitae, publication list, a statement of research interests and arrange for two letters of recommendation. Materials should be sent by January 15th, 1997 to: Dr. Luis A. Aguilar Postdoctoral Committee Instituto de Astronomia/UNAM P.O. Box 439027 San Ysidro, CA 92143-9027 U.S.A. Notifications will be made by February 15th, 1997.

Postdoctoral Position at Lick Observatory

The University of California Observatories/Lick Observatory invites applications for Postgraduate Researcher position at UCO/Lick Observatory. Contingent on the availability of funding, the appointment will be for two years beginning fall 1997 (start date negotiable). The appointment may be extended if funds allow. The appointee will work with Dr. Dennis Zaritsky on the analysis of IPAC, ISO, and HST data in combination with ground-based images and spectra to study the properties of dust in galaxies. There will be an opportunity for the appointee to work on independent research. The appointee will have access to an ongoing UBVI photometric survey of the Magellanic Clouds. Preference will be given to candidates who have worked either on topics related to dust in galaxies, the Magellanic Clouds, or stellar populations.

Applicants should send a vita, a brief (two page limit) summary of research experience and interests related to the topics listed above, and the names of at least three referees (including postal and e-mail addresses, and phone numbers) who have been asked by the applicant to submit letters of reference to Postgraduate Researcher Recruitment c/o UCO/Lick, Office of the Director University of California Santa Cruz, CA 95064

Refer to position #T96-09 in your reply. Closing date for all materials is Jan. 15, 1997 If any questions arise please contact D. Zaritsky at dennis@ucolick.org

Meeting Announcements

Hot Stars in Open Clusters of The Galaxy and Magellanic Clouds

Preliminary Announcement: Workshop Celebrating the 100th Anniversary of La Plata University December 1997, LA PLATA, Argentina

Topics to be discussed:

– Intrinsic properties of hot stars in environments of different metallicities: observations in different spectral domains.

– Models of hot star atmospheres.

– Evolutionary models of hot stars.

– Demography of hot stars in clusters and associations.

– IMF (PDMF), multiplicity, sequential star formation Scientific Organizing Committee: A.Feinstein (Argentina, chair); D.Lennon (Germany); H.Levato (Argentina); A.Maeder (Switzerland); A.F.J.Moffat (Canada); N.Morrell (Argentina); V.Niemela (Argentina); R.Sagar (India); W.Seggewiss (Germany); N.Walborn (USA) Local Organizing Committee: R.Barba, G.Baume, E.Brandi, C.Feinstein, L.Garcia, N.Morrell (co-chair), V.Niemela (co-chair), R.Vazquez. Contact: Postal address: Observatorio Astronómico, Paseo del Bosque, La Plata 1900, Argentina Tel/Fax: 54-21-25-8985

e-mail: hoc97@fcaglp.fcaglp.unlp.edu.ar